OpenStudy (anonymous):

4 underroot x^9y^7/ xy^3 find the radical with exponents?Plzz help me understand this one

OpenStudy (anonymous):



OpenStudy (anonymous):

\[4\sqrt{x^9y^7}/{xy^3}\]

OpenStudy (anonymous):

plz help

OpenStudy (phi):

I assume that is the 4th root (and not just multiplying by 4) you can "pull out" sets of 4. for example \( \sqrt[4]{x*x*x*x} \) becomes x (pull out all 4 x's) x^9 means x times itself 9 times. you can break it up into groups of 4: (x*x*x*x) * (x*x*x*x) *x (you can see why people decided to use exponents) each "quadruple of x's" can be "pulled out" and changed to x so \[ \sqrt[4]{(x*x*x*x) * (x*x*x*x) *x} = x\cdot x\cdot \sqrt[4]{x} \] \[ = x^2\sqrt[4]{x} \] can you do the problem \[ \sqrt[4]{x^9 y^7} \] ? if you can you are ready for \[ \frac{\sqrt[4]{x^9 y^7}}{xy^3} \] The final step is divide by xy^3